# CuckooGraph: A Scalable and Space-Time Efficient Data Structure for Large-Scale Dynamic Graphs

### v1

@@ Abstract
Graphs play an increasingly important role in various big data applications. However, existing graph data structures cannot simultaneously address the performance bottlenecks caused by the dynamic updates, large scale, and high query complexity of current graphs. This paper proposes a novel data structure for large-scale dynamic graphs called CuckooGraph. It does not need to know the amount of graph data in advance, and can adaptively resize to the most memory-efficient form according to the data scale, realizing multiple graph analytic tasks faster. The key techniques of CuckooGraph include Transformation and Denylist. Transformation fully utilizes the limited memory by designing related data structures that allow flexible space transformations to smoothly expand/tighten the required space depending on the number of incoming items. Denylist efficiently handles item insertion failures and further improves processing speed. We conduct extensive experiments, and the results show that CuckooGraph significantly reduces query time by four orders of magnitude on 1-hop successor and precursor queries compared to the state-of-the-art.
@@ I-A Background and Motivation
Graphs can intuitively represent various relationships between entities and are widely used in various big data applications, such as user behavior analysis in social/e-commerce networks [1, 2, 3], financial fraud detection in transactional systems [4, 5, 6], network security and monitoring in the Internet [7, 8, 9], and even trajectory tracking of close contacts of the COVID-19 epidemic [10, 11], $etc.$ Correspondingly, graph analytic systems also play an increasingly significant role, responsible for storing, processing, and analyzing graph-like data well.
@@
As an essential part of graph analytics systems, graph storage schemes are currently facing challenges introduced by the following properties of graph data: ① Fast update: graphs always arrive quickly and are highly dynamic [12, 13, 14]. This requires the storage structure to be updated at high speed. ② Large data scale: graphs can even reach hundreds of millions of edges [15, 16, 17]. This requires the storage structure to be flexibly adapted to the data scale. ③ High query complexity: graphs have complex topologies, and their node degrees are often unevenly distributed and follow a power-law distribution [18, 19, 20]. This means that graphs usually consist of mostly low-degree nodes and a few high-degree nodes. Querying the neighbors of high-degree nodes takes longer, while low-degree nodes take less time. However, the former is more likely to be queried and updated than the latter. This imbalance leads to poor query performance and hinders further optimization. In summary, an ideal graph storage scheme can achieve memory saving, fast running speed, and good update and expansion performance to deal with any unknown graphs.
Unfortunately, neither existing graph storage nor graph database schemes can simultaneously address ①②③ above. Most of them adopt adjacency lists (mainly for sparse graphs) or adjacency matrices (mainly for dense graphs) as the basic data structure to store complete graph information (edges, nodes, $etc.$). However, they are unable to achieve both acceptable space and time efficiency when dealing with large-scale dynamic graphs. Take schemes [21, 22, 23, 24, 25, 26, 27, 28, 29, 30] based on adjacency list (or similar) structures as an example: they certainly cannot address ③. For querying a given edge $\left\langle u,v\right\rangle$, they have to traverse the entire list associated with the starting $u$, so the query speed depends heavily on the degree of the node $u$. Recently, Wind-Bell Index (WBI) [31] is the state-of-the-art (SOTA) solution for addressing ③. Its data structure consists of an adjacency matrix and many adjacency lists, combining their advantages. In addition, it also speeds up queries by introducing hashing. However, it is not designed for scalability and thus is not friendly enough for ②. Besides, it also has a lot of room for improvement in ① because there is still the possibility of traversing more adjacency lists.
Inspired by the strengths and weaknesses of existing schemes, the aim of this paper is to propose a novel hash-based storage structure to simultaneously address ①②③ well. More intuitively, we should answer the question: How to efficiently store large-scale and fast-updating dynamic graphs, as well as query multiple critical graph query tasks faster?
@@ I-B Our Proposed Solution
Towards the design goal presented in §$\S$ I-A, we propose a novel data structure for storing large-scale dynamic graphs, namely CuckooGraph. It has the following advantages: 1) It enables fast and error-free storage of arbitrary graphs (in terms of scale and origin); 2) It can be flexibly expanded and contracted according to arbitrary graphs of unpredictable size; 3) It guarantees very fast query speed for multiple graph query tasks; 4) It has a high loading rate that can be preset.
@@
The design philosophy of CuckooGraph is as follows. To ensure more time- and space-efficient insertion and query of large-scale dynamic graphs, instead of using traditional adjacency matrices or adjacency lists, we choose a (larger) cuckoo hash table (L-CHT) as the basic data structure with a finer-grained division of space in each bucket. Initially, we assume that edge $\left\langle u,v\right\rangle$ is mapped/stored in this bucket. Next, L-CHT decides whether to perform the transformation based on the degree (the number of incoming $v$) of the node $u$: 1) When the node degree is small: this sparsity is more consistent with most graphs in reality, then we sequentially store the incoming $v$ into the designated divided subspace in the bucket. 2) When the node degree exceeds the specified number of subspaces: these subspaces merged in pairs to form several slots with one of them deposited into the first pointer to the first (smaller) cuckoo hash table (S-CHT) that has just been activated, and all $v$ is transferred into that S-CHT to accommodate more incoming $v$. 3) When the node degree is even larger: S-CHT is incremented with some regularity, $i.e.$, by smoothly increasing the memory space to robustly cope with a massive amount of $v$; and of course, it can also be decremented with some regularity to handle $v$ deletions. In short, our transformation technique (also applicable to L-CHT) make good use of limited memory space and adapt to the gradual increase or decrease of $v$ by smoothly expanding or tightening space through a series of spatial transformations.
@@
Although we seem to significantly improve the time and space efficiency through well-designed L/S-CHT, the shortcomings of cuckoo hashing itself have not yet been addressed. In other words, as the memory space becomes tight due to the increase in incoming items, frequent item replacements caused by hash collisions and eventually insertion failures may occur. This results in either our storage scheme not being error-free or becoming less efficient. Therefore, we further propose the denylist technique, aiming to cooperate with the above-mentioned transformation technique to accurately accommodate those items that ultimately fail to be inserted. By sacrificing a very small amount of memory, it efficiently ensures that our solution is error-free and faster. See §$\S$ III for more details on the two key techniques of CuckooGraph. Further, we also provide the weighted version of CuckooGraph customized for streaming scenarios.
@@
The rest of this paper is organized as follows. We theoretically prove that the time and memory cost of CuckooGraph is desirable through mathematical analysis in §$\S$ IV. We conduct comprehensive experiments on large-scale real-world and synthetic datasets to evaluate the performance of CuckooGraph on multiple graph query and analytic tasks, as detailed in §$\S$ V.
Main Experimental Results: 1) For the edge query task, the Average Query Time (AQT) of CuckooGraph is 51.5×51.5$\times$51.5 × faster than that of WBI, while its memory usage is 2.56×2.56$\times$2.56 × less than that of WBI with a loading rate of 0.9; 2) For the 1-hop successor and precursor query tasks, the AQT of CuckooGraph is 13500.8×13500.8$\times$13500.8 × and 12839.8×12839.8$\times$12839.8 × faster than that of WBI, respectively; 3) For the task of finding triangles, the AQT of CuckooGraph is 136.5×136.5$\times$136.5 × faster than that of WBI; 4) For the task of Single-Source Shortest Paths, the Execution Time of CuckooGraph is 132.8×132.8$\times$132.8 × faster than that of WBI.
@@ II Related Work
In this section, we introduce the most basic data structure of CuckooGraph, $i.e.$, cuckoo hash table (CHT), in §$\S$ II-A, and the SOTA graph storage scheme, $i.e.$, Wind-Bell Index (WBI), in §$\S$ II-B, respectively.
II-A Cuckoo Hash Table
The key idea of cuckoo hash table (CHT) [32] is as follows. Each newly inserted item is mapped to two positions by two hash functions, and an empty bucket is selected from these two positions for insertion. If there are items recorded in both positions, CHT selects one of them to kick out, and the kicked out item needs to be re-inserted: $i.e.$, it is inserted into another corresponding position. On the premise that the two hash functions are independent of each other, CHT has a high loading rate, and its query speed still maintains $O(1)$ of ordinary hash tables. However, the disadvantages of CHT are also obvious. When a CHT has a high loading rate, its remaining memory space is tight, which tends to cause each new item insertion often accompanied by a large number of kick-outs and re-insertions. Thus, the time complexity of such insertions is difficult to estimate. When a CHT is full, the insertion of a new item may directly lead to an infinite loop. Therefore, CHT needs to be set with an insertion upper limit, which leads to the evil result that the usual hash table does not exist: insertion failure. In the worst case, its insertion takes a lot of time, and the end result still fails.
II-B Wind-Bell Index
Wind-Bell Index (WBI) [31] is a SOTA data structure for graph database, which consists of an adjacency matrix and many adjacency lists. Each bucket of the adjacency matrix corresponds to a adjacency list through a pointer. The key idea of WBI is to select the shortest hanging lists through multiple hashes, to address the query inefficiency caused by the imbalance of node degrees that has not been considered in existing graph storage/database schemes. In this way, only a small number of memory accesses are required when WBI queries each edge, which speeds up the edge query. Therefore, for a sparse graph with $n$ nodes, WBI can achieve better load balancing by setting the size of the matrix to $\sqrt{n}\times\sqrt{n}$. Conversely, the traditional adjacency matrix requires the size of $n\times n$, and adjacency lists alone are unable to cope with this issue.
@@ IV Mathematical Analysis
In this section, we theoretically analyze the performance of CuckooGraph. Specifically, we show the time and memory complexity of CuckooGraph.
@@ IV-A Time Cost of CuckooGraph
Since the $LR$ is defined as $\frac{n}{dm}$, by setting $\varepsilon=\frac{dm}{n}-1\geq\frac{1}{G}-1$, the expected time cost for inserting an item can be calculated as $(\frac{G}{1-G})^{O(\log d)}$. If we set $T$ as the maximum number of loops for L-CHT, then the worst-case insertion time cost can be further written as $O(T)$, or $O(1)$ if $T$ is not very large. Here, we provide an experiment to verify the above: We expand CuckooGraph starting from the minimum length and insert all the edges of NotreDame Dataset into it in sequence. It can be calculated that the average number of insertions per item in L-CHT and S-CHT considering the expansion is about 1.017 and 1.006, respectively, which is much less than $T$ ($T=250$ in the experiments in Section V).
@@ IV-B Memory Cost of CuckooGraph
TABLE III: Notes for the abbreviated legends of the experiments.
Abbreviations in Legends Related Metrics Explanations
Ours (NT) and Ours (T) Throughput Results of CuckooGraph’s L-CHT without transformation (NT) and with transformation (T), where NT means the size of the dataset is known in advance (so the L-CHT size can be preset) and T means the size of the dataset is not known in advance (so L-CHT has to be transformed automatically).
V Experimental Results
In this section, we evaluate the performance of CuckooGraph through extensive experiments. First, we evaluate how key parameters affect CuckooGraph in §$\S$ V-B. Then, we evaluate the excellent properties of CuckooGraph itself such as high loading rate and scalability in §$\S$ V-C and §$\S$ V-D, respectively. Finally, we compare the performance of CuckooGraph and the SOTA on several important graph tasks in §$\S$ V-E-V-I.

### v2

@@ Abstract
Graphs play an increasingly important role in various big data applications. However, existing graph data structures cannot simultaneously address the performance bottlenecks caused by the dynamic updates, large scale, and high query complexity of current graphs. This paper proposes a novel data structure for large-scale dynamic graphs called CuckooGraph. It does not require any prior knowledge of the upcoming graphs, and can adaptively resize to the most memory-efficient form while requiring few memory accesses for very fast graph data processing. The key techniques of CuckooGraph include Transformation and Denylist. Transformation fully utilizes the limited memory by designing related data structures that allow flexible space transformations to smoothly expand/tighten the required space depending on the number of incoming items. Denylist efficiently handles item insertion failures and further improves processing speed. Our experimental results show that compared with the most competitive solution Spruce, CuckooGraph achieves about $\times3$ higher insertion throughput while requiring only about $68\%$ of the memory space.
@@ I-A Background and Motivation
Graphs can intuitively represent various relationships between entities and are widely used in various big data applications, such as user behavior analysis in social/e-commerce networks [1, 2, 3], financial fraud detection in transactional systems [4, 5, 6], network security and monitoring in the Internet [7, 8, 9], and even trajectory tracking of close contacts of the COVID-19 epidemic [10, 11], $etc.$ Correspondingly, graph analytics systems also play an increasingly significant role, responsible for storing, processing, and analyzing graph-like data well.
@@
As an essential part of graph analytics systems, graph storage schemes are facing challenges introduced mainly by the following properties of graph data: ① Fast update: graphs always arrive quickly and are constantly dynamic [12, 13, 14]. This requires the storage structure to be updated at high speed. ② Large data scale: graphs can even reach hundreds of millions of edges [15, 16, 17]. This requires the storage structure to be flexibly adapted to the data scale. ③ High query complexity: graphs have complex topologies, and their node degrees are often unevenly distributed and follow a power-law distribution [18, 19, 20]. This means that graphs usually consist of mostly low-degree nodes and a few high-degree nodes. Querying the neighbors of high-degree nodes takes longer, while low-degree nodes take less time. However, the former is more likely to be queried and updated than the latter. This imbalance leads to poor query performance and hinders further optimization. In summary, an ideal graph storage scheme can achieve memory-saving, fast processing speed, and good update and expansion performance to deal with any unknown graphs.
Currently, most existing graph storage and database schemes [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] use the following two as basic data structures: adjacency list and compressed sparse row (CSR), but neither of them directly supports large-scale dynamic graphs. The most widely used adjacency list represents node connections intuitively and is easy to edit (such as adding or removing edges). However, due to its non-contiguous memory allocation and inefficiency in accessing non-neighbor edges, this pointer-intensive data structure is prone to significant space-time overhead as the graph size grows. The CSR provides a more compact array-based representation that is more memory efficient and suitable for fast traversal. However, the CSR is inherently static and struggles with updates for dynamic graphs, as its update usually requires completely rebuilding the CSR structure, which is computationally expensive and inefficient. In order to accommodate large-scale dynamic graphs, the data structures of many state-of-the-art graph storage schemes are evolved from adjacency lists or CSRs. Unfortunately, they cannot completely avoid the above-mentioned shortcomings of the adjacency list or CSR itself, so that they cannot solve the above ①②③ at the same time or have various limitations, and there is still room for improvement. For example, Spruce [36], the most competitive solution whose basic data structure is based on adjacency lists, still needs to record many pointers.
@@ I-B Our Proposed Solution
In this paper, we propose a novel data structure for storing large-scale dynamic graphs, namely CuckooGraph [^1] Note that CuckooGraph is essentially an algorithm-level solution that pursues faster insertion and query speed and less memory usage to serve graph-related tasks, different from those system-level solutions.. It has the following advantages: 1) It is memory-saving and can be flexibly expanded or contracted according to actual operations; 2) It basically maintains the fastest running speed in a series of graph analytics tasks; 3) It works for any graph of unknown size without knowing any information about it in advance.
@@
The design philosophy of CuckooGraph is as follows. Instead of using the traditional adjacency list or CSR, we choose a hash-array-based data structure to improve time and space efficiency when handling large-scale dynamic graphs. Specifically, we utilize a (large) cuckoo hash table (L-CHT) [42] as the basic data structure with a finer-grained partitioning of the space in each bucket. We assume that the edge $\left\langle u,v\right\rangle$ is is mapped and will be stored in this bucket. Initially, part of the bucket space is used to store node $u$, and the other part, which is divided into an even number of small slots, is used to store node(s) $v$. Then, L-CHT decides whether to perform our transformation technique based on the degree (the number of incoming $v$) of the $u$: 1) When the node degree is small: this sparsity is more consistent with most graphs in reality, then we sequentially store the incoming $v$ into the small slots. 2) When the node degree exceeds the specified number of small slots: these small slots merged in pairs to form several large slots with one of them deposited into the first pointer to the first (small) cuckoo hash table (S-CHT) that has just been activated, and all $v$ is transferred into that large-capacity S-CHT to accommodate more incoming $v$. 3) When the node degree is even larger: S-CHT is incremented with some regularity, to cope with the large increase of $v$; and of course, it can also be decremented with some regularity to handle $v$ deletions. In short, our transformation smoothly expands or shrinks the space through a series of spatial transformations to adapt to the increase or decrease of $v$, while ensuring that few accesses are required even in the worst case. It greatly reduces the number of pointers and makes full use of limited memory space while ensuring speed.
@@
Although we seem to fully guarantee time and space efficiency through the well-designed L/S-CHT, the shortcomings of cuckoo hashing itself have not yet been addressed. As the memory space becomes tight due to the increase in incoming items, item replacement caused by hash collisions may occur frequently and may cause insertion failures. On the one hand, not handling insertion failures may result in CuckooGraph no longer being error-free, which is unacceptable; on the other hand, we can also address it by directly expanding CuckooGraph every time an insertion failure occurs, but this may make it slower. Therefore, we further propose the denylist optimization, which aims to cooperate with the transformation technique to efficiently accommodate those items that fail to be inserted. Our ablation experiments have verified that this optimization can improve insertion and query throughput with almost no additional memory overhead. For more details on CuckooGraph’s transformation technique and denylist optimization, please refer to §$\S$ III-A1 and §$\S$ III-A2, respectively. Further, we also propose an extended version of CuckooGraph for streaming scenarios to support duplicate edges in §$\S$ III-B.
@@
The rest of this paper is organized as follows. We theoretically prove that the time and memory cost of CuckooGraph is desirable through mathematical analysis in §$\S$ IV. We conduct extensive experiments on 7 large-scale graph datasets with different characteristics to evaluate the performance of CuckooGraph on basic tasks and graph analytics tasks. The results clearly show that CuckooGraph has the fastest speed and the lowest memory overhead on almost all tasks, see §$\S$ V for details.
Main Experimental Results: 1) For basic tasks (§$\S$ V-D), the insertion and query throughput of CuckooGraph is on average 32.66×32.$6$ and 133.62×133.62$\times$133.62 × faster than those of Spruce, respectively, while its memory usage is on average 1.47×1.47$\times$1.47 × less than that of Spruce ($i.e.$, only 68.03% of its); 2) For graph analytics tasks (§$\S$ V-E), the running time of CuckooGraph on 7 typical tasks (Breadth-First Search, Single-Source Shortest Paths, Triangle Counting, Connected Components, PageRank, Betweenness Centrality, and Local Clustering Coefficient) is on average 0.73×0.73$\times$0.73 ×, 168.45×168.45$\times$168.45 ×, 21.33×21.$3$, 1.07×1.07$\times$1.07 ×, 1.03×1.03$\times$1.03 ×, 16.17×16.17$\times$16.17 ×, and 5.80×5.80$\times$5.80 × faster than those of Spruce, respectively.
@@ II Related Work
In this section, we introduce graph storage schemes whose main contribution lies in data structure design and Cuckoo Hash Table (CHT)—the most basic data structure of CuckooGraph, in §$\S$ II-A and §$\S$ II-B, respectively.
II-A Existing Solutions
There are many existing works with the concept of dynamic graph storage, only some of which focus on optimizing graph updates (insertion, deletion, and attribute change, $etc.$) at the algorithm level. Most of them have data structures based on or improved on adjacency lists or CSRs. Of course, there are also other or hybrid ones. We only select representatives to introduce for the sake of space.
Adjacency list-based: GraphOne [29] uses a complementary combination of adjacency lists and edge log lists. The adjacency list stores snapshots of old data, while the edge log records the latest updates, which periodically transfers data into the adjacency list in batches. LiveGraph [30] uses the proposed Transactional Edge Log (TEL) and Vertex Blocks (VB) to store edge information and nodes, respectively. In TEL, edge insertions and updates are performed in the form of log entries in a specified order. RisGraph [31] uses the proposed indexed adjacency lists and sparse arrays. The indexed adjacency list is a dynamic array including arrays and edge indexes for continuous storage, while the sparse array is used for updates to avoid unnecessary data access. Sortledton [34] uses a customized adjacency list, including the expandable adjacency index and adjacency set, to store the mapping from nodes to edge sets and the neighbors of each node, respectively. Wind-Bell Index (WBI) [35] consists of an adjacency matrix and many adjacency lists, where each bucket of the matrix is associated with an adjacency list through a pointer. It selects the shortest hanging list through multiple hashes to address the slow query caused by node degree imbalance that is not considered in existing graph databases. Spruce [36] consists of an edge-storage part and a node-indexing part similar to the vEB tree. The edge-storage part is based on the adjacency list and is used to store the edges as well as attributes. The node-indexing part includes a hash table and a bit vector. It is used to record node identifiers and map nodes to their connected edges. Also, it divides the 8-byte identifier into 4, 2, 2, where 4 is stored in the hash table to share the same hash address, and two 2s are stored in the bit vector associated with the edge storage part. In this way, Spruce achieves low memory consumption and efficient dynamic operations, but it still needs to record quite a few pointers to any graph.
@@ II-A Existing Solutions
CSR-based: To address the problem that CSR is not suitable for dynamic graphs, PCSR [26] replaces the array storing neighbors in CSR with the packed memory array (PMA) [^2] PMA is a dynamic array used to maintain an ordered collection of items. It balances item insertion and deletion operations by interspersing empty slots within the array to optimize access and modification performance. [43], which essentially maintains an implicit complete binary search tree. To avoid frequent rebalance of PCSR, VCSR [33] maintains PMA by reserving empty slots between nodes in proportion to the current node degree. Teseo [44] mainly uses PMAs as leaf nodes of improved B+ trees for graph updates, but only supports undirected graphs. Each PMA is divided into several expandable segments, and a hash table maps nodes to locations within the segment.
Other-based: Terrace [13] decides which data structure to use for storage based on the node degree: 1) Nodes with few neighbors are stored in an sorted array; 2) Nodes with a medium number of neighbors are stored in a PMA; 3) Nodes with many neighbors are stored in a B+ tree. However, its biggest drawback is that the number of nodes for the workload must be known in advance.
II-B Cuckoo Hashing
Cuckoo Hashing (or Cuckoo Hash Table, CHT) [42] contains two hash tables, each associated with a hash function. Each newly inserted item is mapped to two candidate buckets (one in each table), and one of the two buckets will be selected to store it. If at least one of the two candidate buckets is empty, CHT stores the item in an empty bucket. If both candidate buckets are full, CHT randomly selects one of the stored items to kick out, and the kicked out item will be re-inserted into its another candidate bucket. This process is repeated until each item finds a bucket to settle down, or reaches the maximum kick-out threshold and has to exit, which means that there is an item insertion failure. To query an item, CHT only needs to directly check the two candidate buckets through two hash functions. Therefore, CHT has a high loading rate and $O(1)$ query time complexity. However, in the worst case, item insertions take a lot of time while still failing as described above, especially when CHT has a high loading rate.
@@ IV Mathematical Analysis
In this section, we theoretically analyze the performance of CuckooGraph (basic version). Specifically, we show its time and memory complexity.
@@ IV-A Time Cost of CuckooGraph
Since the $LR$ is defined as $\frac{n}{dm}$, by setting $\varepsilon=\frac{dm}{n}-1\geq\frac{1}{G}-1$, the expected time cost for inserting an item can be calculated as $(\frac{G}{1-G})^{O(\log d)}$. If we set $T$ as the maximum number of loops for L-CHT, then the worst-case insertion time cost can be further written as $O(T)$, or $O(1)$ if $T$ is not very large. Here, we provide an experiment to verify the above: We expand CuckooGraph starting from the minimum length and insert all the edges of NotreDame Dataset into it in sequence. It can be calculated that the average number of insertions per item in L-CHT and S-CHT considering the expansion is about 1.017 and 1.006, respectively, which is much less than $T$ ($T=250$ in the experiments in §$\S$ V).
@@ IV-B Memory Cost of CuckooGraph
V Evaluation
In this section, we evaluate the performance of CuckooGraph through extensive experiments, which are briefly described as follows: 1) We introduce the experimental setup in §$\S$ V-A; 2) We evaluate how key parameters affect CuckooGraph in §$\S$ V-B; 3) We verify the effect of Denylist optimization through ablation experiments in §$\S$ V-C; 4) We evaluate the insertion, query, and deletion throughput as well as memory usage of CuckooGraph and its competitors in §$\S$ V-D; 5) We evaluate the running time of CuckooGraph and its competitors on graph analytics tasks (BFS, SSSP, TC, CC, PR, BC, LCC) in §$\S$ V-E.
TABLE III: A brief analysis of the graph datasets used.
Graph Dataset Weighted? # Nodes # Edges # Edges (dedup) Avg. Deg. Max. Deg. Edge Density
CAIDA 0.51M 27.12M 0.85M $1.66$ $17950$ $3.26\times 10^{-6}$
NotreDame 0.33M 1.50M 1.50M $4.60$ $10721$ $1.41\times 10^{-5}$
StackOverflow 2.60M 63.50M 36.23M $13.92$ $60406$ $5.35\times 10^{-6}$
WikiTalk 2.99M 24.98M 9.38M $3.14$ $146311$ $1.05\times 10^{-6}$
Weibo 58.66M 261.32M 261.32M $4.46$ $278491$ $7.60\times 10^{-8}$
DenseGraph 8K 57.59M 57.59M $7199.16$ $14537$ $0.90$
SparseGraph 5M 30M 30M $6$ $6$ $1.20\times 10^{-6}$